# Secret Breach Detection in Source Code with Large Language Models

### v1

@@ Abstract
Source Code, Secret. Large Language Model, Regular Expression
@@ I Introduction
Platforms like GitHub and Bitbucket have become essential for developers to collaborate and share code. Unfortunately, this also means that secrets are often inadvertently exposed within repositories. GitGuardian’s annual report on secret leaks highlights a sharp increase in such incidents since 2020. In 2022 alone, 10 million hardcoded secrets were detected—marking a 67% rise compared to 2021—while scanned commits surged from 860 million in 2021 to over 1 billion in 2022 [3]. Numerous reports have documented cases where attackers exploited exposed API keys from public repositories to access free services [4]. A study by Sophos revealed that compromised credentials were responsible for half of all recorded cyberattacks in the first half of 2023 [5]. Despite efforts by cloud providers like Amazon and Google to detect and revoke leaked keys, version control systems (VCS) offer minimal protections. Beyond source code, secrets may also be unintentionally leaked through configuration files, test scripts, documentation, dependency files, and logs [2].
The 2024 report [6] also identified a significant rise in GenAI-related secret leaks, most notably a 1212 times increase in leaked OpenAI API keys, making them the highest growth detector class. These secrets were leaked at an alarming rate of 46,000 per month on average. Equally concerning is the rise of “zombie leaks”—secrets that remain valid even after repositories are deleted or made private. Such secrets are especially dangerous as they linger without visibility, creating long-term vulnerabilities.
@@
To address these limitations, Meli et al. [15] conducted a comprehensive study on GitHub repositories, combining pattern matching with entropy-based filtering via regex. Their approach successfully detected secrets in over 100,000 repositories with an accuracy of 99.29%. Further advancements were made by Sinha et al. [16], who incorporated keyword searches and program slicing techniques to improve both recall and precision in identifying API keys and similar credentials. Meanwhile, Saha et al. [17] tackled false positives by training machine learning models on datasets retrieved through the Git REST API [18], achieving an F1-score of 86.7% using a Soft-Voting Classifier. Feng et al. [19] further refined secret detection by applying deep neural networks to reduce false positives in textual password detection. To mitigate the risks associated with secret leaks, Sinha et al. [16] suggested various preventive measures, including the use of tools like git-crypt [20], which encrypts sensitive data in Git repositories. Wahab et al. [21] used transformer models like BERT [22] to detect secrets in Github issue reports. Additionally, Basak et al. [23] compiled 24 best practices for effective secret management, offering structured guidelines to minimize exposure risks
In this work, we propose a novel approach to automate secret detection in source code by leveraging large language models (LLMs). Our methodology begins with regex-based candidate string extraction, after which an LLM classifies true secrets, significantly reducing false positives. We experiment with various strategies, including zero-shot, one-shot, few-shot learning, and zero-shot inference after fine-tuning. Our fine-tuned model achieves an F1-score of 0.9852 on the test dataset, outperforming regex-based and conventional machine learning approaches.
@@
The rest of this paper is organized as follows. Section II-A describes the methodology, including data collection, preprocessing, and secret detection using large language models. Section II-B presents our LLM-based detection framework, covering prompting strategies and performance across different models. Section II-C discusses the use of fine-tuned smaller LLMs for local deployment. Section III provides a detailed discussion on evaluation insights, limitations, and impact factors. Section IV reviews related work. Section V outlines potential threats to validity, and finally, Section VI concludes the paper and discusses future directions.
@@
II Methodology
II-A Data Collection and Preprocesssing
II-A1 Data Collection
@@ II-A2 Preprocessing
For multiclass classification, we only took the positive samples from the dataset. The dataset distribution is mentioned in table I.
@@
II-B Secret Detection Using Pre-trained Large Language Models (RQ1)
II-B1 Motivation
Pre-trained large language models (LLMs) have shown remarkable performance have demonstrated exceptional capabilities in text classification [26] and contextual understanding [27], making them well-suited for secret detection. Their ability to understand code semantics [28] without extensive task-specific fine-tuning makes them attractive for security-related applications as well. Motivated by this, we investigate the inherent capability of LLMs to identify secrets embedded in source code. Specifically, we explore how well these models can distinguish true secrets from non-sensitive strings using contextual understanding, even without extensive retraining.
II-B2 Approach
@@
Evaluation Data: For evaluation, we randomly selected a subset of 3000 (1500 secret + 1500 non-secret) data points from our dataset.
Evaluation Metrics: To assess the performance of our LLM-based secret detection models, we employ standard classification evaluation metrics: precision, recall, and F1-score. These metrics provide insights into how well the model distinguishes between secrets and non-sensitive strings.
@@
  • F1-Score: The F1-score is the harmonic mean of precision and recall. It provides a balanced performance measure, especially useful when both false positives and false negatives are important to minimize.
Justification for Using F1-Score Since our dataset is balanced (12,000 secrets and 12,000 non-secrets), accuracy serves as a reasonable metric during training. However, in real-world scenarios, secret detection often requires balancing precision and recall. A model focused too heavily on precision may miss actual secrets, while one optimized for recall may trigger too many false alarms. To address this, we use the F1-score as our primary evaluation metric. It provides a balanced measure that minimizes both false positives and false negatives, which is crucial in security workflows where over-detection causes alert fatigue and under-detection can lead to serious breaches.
TABLE II: Evaluation Metrics for zero-shot learning on pretrained llms
Model Precision Recall F1-Score
DeepSeek-V3 0.7810 0.9033 0.8377
Gpt-4o 0.8233 0.8633 0.8428
TABLE III: Evaluation Metrics for one-shot learning on pretrained llms
Model Precision Recall F1-Score
DeepSeek-V3 0.7761 0.9913 0.8716
Gpt-4o 0.8623 0.9813 0.9180
TABLE IV: Evaluation Metrics for few-shot learning on pretrained llms
Model Precision Recall F1-Score
DeepSeek-V3 0.7926 0.9887 0.8799
Gpt-4o 0.9329 0.9465 0.9392
II-B3 Result
The results of DeepSeek and Gpt-4o, as presented in Table II, Table III, and Table IV, demonstrate that Gpt-4o consistently outperforms DeepSeek across all evaluation metrics—Precision, Recall, and F1-Score—in zero-shot, one-shot, and few-shot settings.
  • Zero-Shot: In the zero-shot setting, Gpt-4o achieves a Precision of 82.33%, Recall of 86.33%, and an F1-Score of 84.28%, outperforming DeepSeek across all metrics. This suggests that Gpt-4o is better equipped to utilize its pre-trained knowledge to identify sensitive strings from code snippets even without any labeled examples.
  • One-Shot: With the addition of a single labeled example, Gpt-4o shows substantial improvement—particularly in Precision, which rises to 86.23%, and F1-Score, which reaches 91.80%. Compared to DeepSeek, Gpt-4o demonstrates stronger generalization from minimal guidance.
  • Few-Shot: In the few-shot setting, Gpt-4o achieves its highest scores: 93.29% Precision, 94.65% Recall, and 93.92% F1-Score. These results surpass those of DeepSeek and indicate that Gpt-4o can effectively leverage multiple examples to distinguish secrets from non-sensitive content with exceptional accuracy and coverage.
II-B4 Multiclass Classification
As mentioned in Section II-A we extended our approach to multiclass classification. We evaluated using 3000 test samples from the partition mentioned in the dataset. Despite their general language understanding capabilities, these models exhibit significantly lower performance in secret type classification compared to our finetuned counterparts.
Table V presents the overall classification performance, with GPT-4o achieving a weighted F1-score of 0.8108. The weighted F1-score was chosen as the evaluation metric to account for class imbalance in the dataset.
TABLE V: Overall Multiclass Classification Accuracy for Pretrained Models
Model F1-score(weighted)
DeepSeek-V3 0.7338
GPT-4o 0.8108
Table VI presents the detailed classification report. While classes like Private Key and API Key and Secret achieve relatively high precision and recall, other categories such as Generic Secret, Other, and Username show poor performance, suggesting that zero-shot or few-shot capabilities are insufficient for nuanced secret-type recognition.
TABLE VI: Multiclass Classification Report for GPT-4o
Category Precision Recall F1-Score
Private Key 0.9897 0.9411 0.9648
API Key and Secret 0.7491 0.6667 0.7055
Authentication Key and Token 0.9009 0.8276 0.8627
Other 0.2373 0.1429 0.1783
Generic Secret 0.0844 0.4407 0.1417
Database and Server URL 0.5581 0.6857 0.6154
Password 0.8485 0.7179 0.7778
Username 0.0000 0.0000 0.0000
Weighted F1-score 0.8108
RQ1 Summary: This study investigates the ability of pre-trained large language models (LLMs), specifically GPT-4o and DeepSeek-V3, to detect secrets in source code without extensive fine-tuning. By using context windows of 200 characters and structured prompts, the models were evaluated in zero-shot, one-shot, and few-shot learning settings. GPT-4o consistently outperformed DeepSeek across all prompting strategies, achieving the highest F1-score of 93.92% in the few-shot setting. Although effective at binary classification, the models did not perform remarkably well in multiclass secret-type classification, highlighting the need for fine-tuning to achieve reliable performance in nuanced tasks.
II-C Secret Detection with Smaller Fine-tuned Models (RQ2)
II-C1 Motivation
@@ II-C2 Approach
Parameter Settings: For fine-tuning our models, we employed QLoRA [38] with 4-bit precision to minimize memory consumption while preserving model effectiveness. Specifically, we used NF4 (Normal Float 4-bit) quantization with FP16 computation, and disabled double quantization to streamline training. Fine-tuning was carried out on transformer attention layers, enabling the models to learn domain-specific patterns crucial for accurate secret classification, even on resource-constrained hardware.
We adopted parameter-efficient fine-tuning (PEFT) [39] using the LoRA framework, which allows selective updating of model parameters. We set the LoRA rank to 64 and scaling factor $\alpha$ to 16. No dropout or bias modifications were used, ensuring faster training and reduced overhead. This setup facilitates robust performance while requiring only a fraction of the resources compared to full fine-tuning.
The models were fine-tuned using the Paged AdamW (32-bit) optimizer with a learning rate of 2e-4. This optimizer was chosen for its memory-efficient paging mechanism, which enables training large models on limited GPU resources without compromising performance. Training was performed for 7 epochs with a batch size of 1 per device and 8 gradient accumulation steps. A cosine decay learning rate scheduler was used, with a warm-up ratio of 3%.
II-C3 Evaluation Data
@@
  • Test Set: 1,500 true secrets and 1,500 non-secrets (3,000 total, same set as Section II-B)
@@
Evaluation Metrics: Similar to Section II-B, we used those evaluation metrics to evaluate the detected secrets.
@@
TABLE VII: Evaluation Metrics for Zero-Shot Prompting (Raw Models)
Model Precision Recall F1-Score
DeepSeek-7B 0.4105 0.4283 0.4054
Gemma-7B 0.5501 0.5003 0.3353
LLaMA-3.1 8B 0.5412 0.5130 0.4125
Mistral-7B 0.5033 0.5033 0.5029
CodeLLaMA-7B 0.5453 0.5383 0.5207
TABLE VIII: Evaluation Metrics for One-Shot Prompting (Raw Models)
Model Precision Recall F1-Score
DeepSeek-7B 0.6864 0.5743 0.4990
Gemma-7B 0.6741 0.6610 0.6545
LLaMA-3.1-8B 0.7622 0.6920 0.6699
Mistral-7B 0.6848 0.5720 0.4949
CodeLLaMA-7B 0.6741 0.6610 0.6545
TABLE IX: Evaluation Metrics for Few-Shot Prompting (Raw Models)
Model Precision Recall F1-Score
DeepSeek-7B 0.5216 0.5127 0.4563
Gemma-7B 0.7286 0.7187 0.7156
LLaMA-3.1-8B 0.7913 0.7913 0.7913
Mistral-7B 0.7678 0.7630 0.7619
CodeLLaMA-7B 0.7286 0.7187 0.7156

### v2

@@ Abstract
Source Code, Secret, Large Language Model, Regular Expression
publicationid: pubid: 979-8-3315-9147-2/25/\$31.00 ©2025 IEEE
@@ I Introduction
Platforms like GitHub and Bitbucket are vital for developer collaboration but often lead to inadvertent secret leaks. GitGuardian’s reports show a sharp rise in such incidents: in 2022 alone, 10 million hardcoded secrets were detected—a 67% increase from 2021—as scanned commits rose from 860 million to over 1 billion [3]. Exposed API keys have enabled attackers to exploit free services [4], and Sophos found that compromised credentials caused half of all cyberattacks in early 2023 [5]. Despite detection efforts by cloud providers, version control systems provide limited protection. Leaks occur not only in source code but also in config files, test scripts, docs, dependency files, and logs [2]. The 2024 report [6] highlighted a 1212× surge in leaked OpenAI API keys—the fastest-growing secret type—at a rate of 46,000 per month. Particularly alarming are “zombie leaks,” which remain valid even after repos are deleted or privatized, posing long-term risks.
@@
To address these limitations, Meli et al. [15] conducted a comprehensive study on GitHub repositories, combining pattern matching with entropy-based filtering via regex. Their approach successfully detected secrets in over 100,000 repositories with an accuracy of 99.29%. Further advancements were made by Sinha et al. [16], who incorporated keyword searches and program slicing techniques to improve both recall and precision in identifying API keys and similar credentials. Meanwhile, Saha et al. [17] tackled false positives by training machine learning models on datasets retrieved through the Git REST API [18], achieving an F1-score of 86.7% using a Soft-Voting Classifier. Feng et al. [19] further refined secret detection by applying deep neural networks to reduce false positives in textual password detection. To mitigate the risks associated with secret leaks, Sinha et al. [16] suggested various preventive measures, including the use of tools like git-crypt [20], which encrypts sensitive data in Git repositories. Wahab et al. [21] used transformer models like BERT [22] to detect secrets in Github issue reports.
In this work, we propose a novel approach to automate secret detection in source code by leveraging large language models (LLMs). Our methodology begins with regex-based candidate string extraction, after which an LLM classifies true secrets, significantly reducing false positives. We explore several strategies, including zero-shot inference both before and after fine-tuning, and one-shot and few-shot prompting on pre-trained models.. Our fine-tuned model achieves an F1-score of 0.9852 on the test dataset, outperforming regex-based and conventional machine learning approaches.
@@
The rest of this paper is organized as follows. Section II reviews related work. Section III-A describes the methodology, including data collection, preprocessing and data distribution. Section III-B presents our LLM-based detection framework, covering prompting strategies and performance across different models. Section III-C discusses the use of fine-tuned smaller LLMs for local deployment. Section V provides a detailed discussion on evaluation insights, limitations, and impact factors. Section VI outlines potential threats to validity, and finally, Section VII concludes the paper and discusses future directions.
@@
II Related Work
Large-Scale Secret Leak Analysis
Meli et al. [15] conducted a large-scale study on secret leaks in GitHub by scanning billions of files through real-time and snapshot-based analysis. Their regex and entropy-based filtering approach achieved high precision (99.29%) but relied on a relatively small set of patterns and minimal manual labeling, potentially limiting coverage and generalizability. In contrast, our work uses a curated benchmark dataset with 768 regex patterns and 15,000 manually labeled secrets, leveraging large language models (LLMs) to incorporate contextual understanding.
Program Analysis for Secret Detection
Building upon earlier approaches, Sinha et al. [16] enhanced secret detection by incorporating program slicing, keyword-based, and pattern-based searches. The approach achieved up to 100% precision and 84% recall for Amazon AWS keys by leveraging static analysis techniques. While impactful, the method is useful for specific patterns and language constructs, limiting its adaptability across diverse codebases.
Secret Detection Tools
Tools like TruffleHog and Gitleaks scan repositories for hardcoded secrets using regexes, entropy checks, and known patterns. While effective for common cases, they often struggle with false positives and lack contextual understanding, especially for obfuscated or non-standard secrets.
Machine Learning-Based Filtering
Saha et al. [17] used regex-extracted data and trained a Soft Voting Classifier on binary features like entropy and character patterns, achieving an F1-score of 86.7%. However, their reliance on handcrafted features and lack of contextual understanding limited their performance. We address this by leveraging LLMs that capture semantic and structural cues from code, yielding significantly higher accuracy.
Deep Learning Approaches
Feng et al. [19] applied deep neural networks to detect textual passwords, reducing false positives through semantic learning. Yet, their focus was narrow and lacked generalization across diverse secret types. Our LLM-based method extends deep learning to diverse secrets in code, supporting both binary and multiclass classification.
Secret Detection in Issue Reports
Comparative Analysis of Detection Tools
III Methodology
III-A Data Collection and Preprocessing
III-A1 Data Collection
@@ III-A2 Preprocessing
For multiclass classification, we only took the positive samples from the dataset. The dataset distribution is mentioned in table I. The Other category includes various forms of sensitive values that do not fall under standard API keys or credentials. These may include a Package Key ID (e.g., PKG-8736-XYZ), an Instrumentation Key (e.g., a1b2c3d4-e5f6-4711-9a8b-112233445566), or a random string used for internal tracking or authentication purposes.
@@
III-B Secret Detection Using Pre-trained Large Language Models (RQ1)
III-B1 Motivation
Pre-trained large language models (LLMs) have shown remarkable performance have demonstrated exceptional capabilities in text classification [25] and contextual understanding [26], making them well-suited for secret detection. Their ability to capture contexts in code [27], even without extensive task-specific fine-tuning, makes them promising for security-related applications. Motivated by this, we investigate the inherent capability of LLMs to identify secrets embedded in source code. Specifically, we explore how well these models can distinguish true secrets from non-sensitive strings using contextual understanding, even without extensive retraining.
III-B2 Approach
@@
Evaluation Data: For evaluation, we randomly selected a subset of 3,000 (1,500 secret + 1,500 non-secret) data points from our dataset.
III-B3 Multiclass Classification
As mentioned in Section III-A we extended our approach to multiclass classification. We evaluated using 3,000 test samples from the partition mentioned in the dataset.
Evaluation Metrics: To assess the performance of our LLM-based secret detection models, we employ standard classification evaluation metrics: precision, recall, F1-score, and F2-score. These metrics provide insights into how well the model distinguishes between secrets and non-sensitive strings.
@@ III-B3 Multiclass Classification
  • F1-Score: The F1-score is the harmonic mean of precision and recall. It provides a balanced performance measure, especially useful when both false positives and false negatives need to be minimized.
  • F2-Score: The F2-score is a weighted harmonic mean of precision and recall, placing more emphasis on recall. This is particularly relevant in security-sensitive applications, where failing to detect a secret (false negative) can have more serious consequences than a false positive.
    $F_{2}=\frac{(1+2^{2})\cdot\text{Precision}\cdot\text{Recall}}{2^{2}\cdot\text{Precision}+\text{Recall}}$ (1)
F1 and F2 Scores for Balancing Precision and Recall: While our dataset is balanced (12,000 secrets and 12,000 non-secrets), real-world secret detection tasks often prioritize capturing as many true secrets as possible. Accuracy is a reasonable metric during training, but it fails to reflect the trade-offs between false positives and false negatives. To evaluate models more effectively, we adopt both the F1-score and F2-score. The F1-score provides a balanced view, minimizing both types of error, while the F2-score emphasizes recall to reduce the likelihood of missed secrets. This dual-metric approach ensures that our evaluation aligns with the practical needs of security workflows—where over-detection can cause alert fatigue, and under-detection can lead to significant data breaches.
III-C Secret Detection with Smaller Fine-tuned Models (RQ2)
III-C1 Motivation
@@ III-C2 Approach
Parameter Settings: For fine-tuning, we used QLoRA [37] with 4-bit precision (NF4 quantization and FP16 computation), disabling double quantization to reduce memory usage. Fine-tuning focused on transformer attention layers for domain-specific pattern learning. We employed parameter-efficient fine-tuning (PEFT) [38] using LoRA with rank 64 and $\alpha=16$, without dropout or bias adjustments. The Paged AdamW optimizer (learning rate $2\times 10^{-4}$) was used with memory-efficient paging. Training ran for 7 epochs with batch size 1, 8 gradient accumulation steps, and a cosine decay scheduler with 3% warm-up.
III-C3 Evaluation Data
@@
  • Test Set: 1,500 true secrets and 1,500 non-secrets (3,000 total, same set as Section III-B)
@@
Evaluation Metrics: Similar to Section III-B, we used those evaluation metrics to evaluate the detected secrets.
III-C4 Multiclass classification
To overcome the limitations of pretrained LLMs, we fine-tuned smaller language models on 9,000 labeled positive samples, specifically curated for the secret classification task. These models were then evaluated on the same 3,000-sample test set used for the pretrained models to ensure a fair comparison.
IV Result
This section presents the evaluation results for our two primary research questions. First, we assess the performance of pre-trained large language models (LLMs), such as GPT-4o and DeepSeek-V3, using various prompting strategies. Then, we investigate the effectiveness of smaller, open-source models that have been fine-tuned for secret detection.
IV-A Secret Detection Using Pre-trained Large Language Models (RQ1)
To answer RQ1, we evaluate GPT-4o and DeepSeek-V3 using zero-, one-, and few-shot prompting to assess their ability to detect secrets without fine-tuning. Table II shows binary classification results, while Tables III and  IV present multiclass performance using precision, recall, F1, and F2 scores.
@@ IV-A Secret Detection Using Pre-trained Large Language Models (RQ1)
TABLE II: Evaluation Metrics for Zero-shot, One-shot, and Few-shot Learning on Pretrained LLMs
Setting Model Precision Recall F1-Score F2-score
Zero-Shot DeepSeek-V3 0.7810 0.9033 0.8377 0.8759
Gpt-4o 0.8233 0.8633 0.8428 0.8550
One-Shot DeepSeek-V3 0.7761 0.9913 0.8716 0.9392
Gpt-4o 0.8623 0.9813 0.9180 0.9549
Few-Shot DeepSeek-V3 0.7926 0.9887 0.8799 0.9421
Gpt-4o 0.9329 0.9465 0.9392 0.9437
The results of DeepSeek-V3 and Gpt-4o, as presented in Table II, show that Gpt-4o consistently outperforms DeepSeek in overall Precision, F1-Score, and F2-Score across zero-shot, one-shot, and few-shot settings. Notably, DeepSeek demonstrates marginally higher Recall in all three settings, highlighting its stronger sensitivity to identifying true positives, albeit at the cost of lower precision.
  • Zero-Shot: DeepSeek-V3 achieves a Recall of 90.33% and a higher F2-Score of 87.59%, compared to Gpt-4o’s Recall of 86.33% and F2-Score of 85.50%. Although Gpt-4o maintains better Precision (82.33% vs 78.10%) and F1-Score (84.28% vs 83.77%), DeepSeek’s superior Recall gives it the edge in F2-Score.
  • One-Shot: With the addition of a single labeled example, Gpt-4o reaches 86.23% Precision, 98.13% Recall, 91.80% F1-Score, and a strong 95.49% F2-Score. Although DeepSeek maintains a slightly higher Recall of 99.13%, its lower Precision results in a reduced F1 (87.16%) and F2 (93.92%) score compared to Gpt-4o.
  • Few-Shot: Gpt-4o attains its best performance with 93.29% Precision, 94.65% Recall, 93.92% F1-Score, and 94.37% F2-Score. DeepSeek again shows slightly higher Recall (98.87%), but falls short in F1 (87.99%) and F2 (94.21%) scores. These results confirm Gpt-4o’s superior ability to generalize effectively when provided with multiple examples.
Table III shows GPT-4o’s classification performance, with a weighted F1-score of 0.8108 and F2-Score of 0.7941. The weighted F1and F2-score was chosen as the evaluation metric to account for class imbalance in the dataset.
TABLE III: Overall Multiclass Classification Accuracy for Pretrained Models
Model F1-score(weighted) F2-score(weighted)
DeepSeek-V3 0.7338 0.7409
GPT-4o 0.8108 0.7941
Table IV presents the detailed classification report. While classes like Private Key and API Key and Secret achieve relatively high precision and recall, other categories such as Generic Secret, Other, and Username show poor performance, suggesting that zero-shot or few-shot capabilities are insufficient for nuanced secret-type recognition.
TABLE IV: Multiclass Classification Report for GPT-4o
Category Precision Recall F1-Score F2-Score
Private Key 0.9897 0.9411 0.9648 0.9504
API Key and Secret 0.7491 0.6667 0.7055 0.6817
Authentication Key and Token 0.9009 0.8276 0.8627 0.8413
Other 0.2373 0.1429 0.1783 0.1553
Generic Secret 0.0844 0.4407 0.1417 0.2390
Database and Server URL 0.5581 0.6857 0.6154 0.6557
Password 0.8485 0.7179 0.7778 0.7407
Username 0.0000 0.0000 0.0000 0.0000
Weighted F1-score 0.8108
Weighted F2-score 0.7941
RQ1 Summary: This study evaluates the ability of pre-trained large language models (LLMs), GPT-4o and DeepSeek-V3, to detect secrets in source code without fine-tuning. Using 200-character context windows and structured prompts, models were tested in zero-shot, one-shot, and few-shot settings. GPT-4o consistently outperformed DeepSeek-V3 in Precision and F1-Score, achieving the highest overall F1-score of 93.92% and F2-score of 94.37% in the few-shot setting. However, DeepSeek-V3 had higher Recall in all cases and a superior F2-score in the zero-shot scenario. While effective for binary classification, both models struggled with multiclass secret-type classification, highlighting the need for fine-tuning for more nuanced tasks.
IV-B Secret Detection Using Small Large Language Models (RQ2)
For RQ2, we fine-tune smaller models (LLaMA-3.1, Mistral, CodeLLaMA, Gemma, DeepSeek) on the SecretBench dataset and compare them to their raw versions. Table V onward summarizes performance, showing that fine-tuning greatly boosts accuracy, even in low-resource settings.